# Geometrical properties of strained and twisted moiré heterostructures

### v1

@@ III.6 Lattice relaxation
The geometrical description in the previous sections assumes that the lattices deform following the externally applied twist and strain. Real systems, however, tend to naturally relax in response to external perturbations, in order to minimize the energy. Relaxation effects in moiré heterostructures have been extensively studied in the literature [van2015relaxation, nam2017lattice, zhang2018structural, carr2018relaxation, guinea2019continuum, lucignano2019crucial, cantele2020structural, leconte2022relaxation, ezzi2024analytical, kang2025analytical]. In the case of twisted bilayer graphene, the relaxation is particularly significant at very small (marginal) twist angles [van2015relaxation], whereby the energy-costly AA stacking regimes shrink and corrugate out of plane, leading to the formation of triangular AB/BA stacking separated by domain walls (DW) [zhang2018structural]. As a result, the relaxed moiré pattern of slightly twisted bilayer graphene can look markedly different than those expected for a rigid configuration. The relaxation acts, by itself, as a source of intrinsic inhomogeneous strain in the system [kazmierczak2021strain, kang2023pseudomagnetic, ceferino2024pseudomagnetic].
@@
Similar relaxation effects are observed under external strain, although to date only few works have studied it. In the particular case of triangular moiré patterns induced by solely biaxial or shear strain (Section III.4.3), Cazeaux et al. reported also an expansion and contraction of AB/BA and AA stacking regimes, respectively, as the strain magnitude decreases [cazeaux2023relaxation]. A similar behavior was also recently reported by Kundu et al. in biaxially strained $\mathrm{MoS_{2}}$ bilayer [kundu2025atomic]. There, the relaxed in-plane strain distribution is also modified by reducing and increasing, respectively, the strain magnitudes at the AB/BA and AA stacking regimes, which as in twisted systems, are separated by domain walls. Since these relaxation effects arise due to local minimization of the stacking and elastic energies, a similar behavior of DW formation is generally expected for non-triangular strained patterns. For instance, the quasi-unidimensional channels (Section III.4.1) also undergo a shrink and expansion of the AA and AB/BA stackings [cazeaux2023relaxation]. Another interesting relaxation effect is the formation of giant atomic swirls under pure biaxial heterostrain [mesple2023giant], whereby the soliton DW swirl around the AA stacking regimes [lalmi2014flower].
@@
In general, the relaxed moiré configurations cannot be accounted by homogeneous twist and strain profiles, as considered in previous sections. Rather, they are characterized by a position dependent displacement vector at each atomic site, which minimizes the elastic and interlayer adhesion energies. The displacement vectors are typically obtained numerically [nam2017lattice, carr2018relaxation, cazeaux2020energy], or analytically [vafek2023continuum, ezzi2024analytical, kang2025analytical]. The geometrical description of relaxed moiré patterns arising from inhomogeneous strain is beyond the scope of this review.
@@
Figure 10: The substrate bending technique. Schematic illustration of generating heterostrain in TBG on a flexible substrate (a) before and (b) after bending. R is the bending radius. (c) Calculated (red line) strain on the flexible substrate surface and experimental (black squares) uniaxial strain in bottom graphene. The thickness of the flexible substrate is 135 $\mathbf{\mu}$m. Inset shows the flexibility of the substrate with TBG device on top. Adapted under the terms of the $C$ BY license from Ref. [gao2021heterostrain]. Copyright (2021) The author(s).
@@ IV Experimental implementation of strain in moiré materials
Due to the nature of nanoscale dimensions and stability, strain is a powerful tuning knob to modify the geometrical and electronic properties of 2D systems [lee2013high, li2015optoelectronic, zhang2018strain, so2021polarization, darlington2020imaging, levy2010strain, bunch2008impermeable, lloyd2016band, thai2021mos2, conley2013bandgap, datye2022strain, castellanos2013local, deng2016wrinkled, lee2020switchable]. In particular, as we discuss in Sec. III, due to the presence of a moiré that acts as a magnifying glass, even very small strains can significantly modify the geometrical and electronic properties of the moiré patterns. In practice, small strain tend to be induced inevitably without control during the sample growth or fabrication, which we refer to as unintentional strain. Intentional strain may be also induced externally and designed carefully by well-established experimental techniques. In this section, we introduce three efficient experimental techniques developed to engineer strain in moiré materials. Then we present various strain-induced moiré pattern geometries generated in experiments.
@@ IV.1 Strain techniques
A large amount of experimental techniques have been introduced to generate strain in 2D materials. For example, strain can be generated by using different microactuators [perez2014controlled, ziss2017comparison], by applying hydrostatic pressure [antoniazzi2024pressure], by placing 2D materials on periodic structures constructed by nanopillars [li2015optoelectronic, jiang2017visualizing], by stacking on piezoelectric stacks [liu2024continuously], or by bending a flexible substrate [wang2019situ]. Here, we will mainly focus on several common approaches that are well-implemented in moiré materials. Note that the strain techniques developed for moiré materials may not be compatible for moiré material-based devices. The strain approaches for 2D material-based devices are out of the scope of this review.
@@ IV.1.1 Substrate out-of-plane bending
Figure 11: Process-induced strain technique. (a) Schematic illustration of generating biaxial heterostrain by fully encapsulating a TBG with a highly stressed thin film. (b) Moiré patterns variation with biaxial heterostrain strength, where $\theta$ is the twist angle in the TBG. (c) Schematic illustration of generating uniaxial heterostrain by partially encapsulating TBG with a highly stressed thin film patterned into a stripe. (d) Moiré patterns variation with uniaxial heterostrain strength, and strain tunable Hubbard model parameters. The Hubbard model parameters $U$ and $t$ are the Coulomb repulsion and hopping terms, respectively. In the process-induced strain approach, the strain magnitude is directly proportional to the applied thin force on the stressor layer. Reprinted from Ref. [pena2023moire], with the permission of AIP Publishing.
@@
Figure 12: Sliding-based strain technique. (a)-(c) Moiré patterns generated by hexagonal lattices with a lattice mismatch $\delta=10\%$ and (a) no strain, (b) $\varepsilon=2$% uniaxial heterostrain, (c) $\varepsilon=10$% uniaxial heterostrain. (d) Slide view schematic of a strained open-face graphene-hBN device. The graphene is stretched by pushing the right electrode with an AFM tip (gray triangle). Red lines show regions where the metal is strongly adhered through the O2–plasma treatment before deposition. (e) Optical image of a device. Red dashed regions are the O2–plasma treated electrode. Purple box labels the graphene channel. Scale bar is 4 $\mathbf{\mu}$m. (f)-(h) 100 nm $\times$ 100 nm CAFM images of graphene-hBN moiré patterns. The $x$-axis of the CAFM is along the graphene channel, parallel to the stretching direction. Insets show FFT of the images. Fitting the wavevectors results in the strain values $\varepsilon$ with strain angles $\phi$ relative to the +x+x-axis, and twist angle $\theta$ between graphene and hBN as (0.04%, +80 $\pm$ 20, -0.12), (0.34%, +1 $\pm$ 5, 0.02) and (0.04%, +17 $\pm$ 4, 0.02) for (f), (g) and (h), respectively. Reprinted with permission from Ref. [sequeira2024manipulating]. Copyright 2024 American Chemical Society.
@@
One strain technique for moiré materials is based on in-plane or out-of-plane stretching or bending the underlying substrate or active 2D materials [wang2019situ, gao2021heterostrain, kapfer2023programming, ou2025continuous]. In this technique, the active 2D layer can be suspended or directly stacked on the substrate. We refer to this method as substrate bending technique. A similar method to generate strain, through in-plane bending of one active 2D layer by the tip of an atomic force microscope (AFM), will be detailed in the following subsection [kapfer2023programming].
@@
Figure 10 presents a generalized approach of dynamically tuning heterostrain in TBG via the substrate out-of-plane bending [gao2021heterostrain]. The setup consists of TBG assembled on a flexible polymethyl methacrylate (PMMA)-coated polyethylene terephthalate (PET) substrate. The twist angle between the two graphene layers is large (about $\theta\sim 13.2^{\circ}$ in the experiment) in order to minimize the interlayer friction force in TBG. As a result, the bottom graphene layer is separated from both the top graphene and the flexible substrate, and the strain generated from the substrate is only transferred to the bottom graphene. Upon bending the PET substrate, because the bottom graphene layer is strongly adhered to the substrate, it becomes uniaxially stretched. The top graphene experiences a negligible amount of deformation because it is bonded by the weak van der Waals force. Consequently, a uniaxial heterostrain is generated in the TBG moiré pattern. The strain strength can be dynamically and in situ modulated by changing the bending radius $R$, following the relation $\epsilon\approx t/2R$, where $t$ is the substrate thickness (Fig. 10(b)). The maximum strain strength achieved by out-of-plane bending in TBG with $\theta\sim 13.2^{\circ}$ was $\sim 1.3\%$ (Fig. 10(c)). The formation of heterostrain can be confirmed by the Raman spectroscopy, which presents triple G peaks with one degenerate mode G peak from unstrained graphene, and two split G+ and G- peaks from strained graphene.
@@ IV.1.2 Process-induced strain
Another common approach is a process-induced strain technique, which is based on the deposition of stressed thin films onto the active 2D materials [pena2021strain, pena2023moire, zhang2024enhancing, zhang2024patternable, liu2025programmable]. In Ref. [pena2023moire], Peña et al. utilized this method to generate designable layer-by-layer strains, on the device-level, with full control over the type of strain (uniaxial or biaxial, tensile or compressive) in moiré systems. By choosing optically transparent stressors, the intentionally induced strain could be characterized through Raman spectroscopy.
@@
Figure 13: Continuous tuning of twist angle and strain in a graphene–hBN heterostructure. (a) Schematic of the continuous tuning approach. The twist angle and the strain are manipulated by bending a 2D material ribbon through a nanomanipulator and the AFM tip. (b) Sketch of a distorted moiré pattern, magnifying the effect of strain in the low-angle limit. (c) (Left panel) AFM images of a bMLG on hBN before (top) and after (bottom) bending, (middle panel) PFM scans taken at the marked spots (white cross and dot) indicated in the bMLG and (right panel) their corresponding FFTs. From Ref. [kapfer2023programming]. Reprinted with permission from AAAS.
@@ IV.1.3 Sliding-based strain
In Ref. [sequeira2024manipulating], Sequeira et al. presented a sliding-based technique to manipulate the uniaxial heterostrain in graphene-hexagonal boron nitride (hBN) and study the strain effect on moiré geometry. To this end, they fabricated an open-faced graphene device on a hBN substrate, with a fixed source electrode and a movable drain electrode. In the electrode region, a metal handle is strongly adhered to the graphene layer (red line in Fig. 12(d)). Consequently, the drain electrode can be moved laterally by pushing it with an AFM tip, inducing heterostrain in the graphene layer. To selectively adhere the electrode to only the graphene layer, a light O2-plasma treatment is applied before depositing the Cr/Au handles. The electrodes are used to both selectively strain and measure transport (Fig. 12(e)). Importantly, the movable electrode has to have sufficient sliding friction to retain the strain in graphene when the AFM tip is retracted. As a result, an open-face graphene sample can be progressively strained, independently of the hBN substrate. The strain effect was measured by performing the conductive AFM (CAFM) measurement on the open graphene channels (Figs. 12(f)-(h)). As seen, the initial graphene aligned with hBN, forming a moiré pattern of length 14 nm, evolves into elongated moiré patterns by stretching the graphene with the AFM tip.
Recently, a similar sliding-based strain method was developed by Carrasco et al. [carrasco2025]. During moiré sample growth, due to the different thermal expansion between graphene and the underlying substrate, graphene layers experience different sources of strain, resulting in graphene wrinkles in the graphene moiré sample. By mechanically manipulating the graphene wrinkles with a scanning tunneling microscopy (STM) tip, they actively tune local strain in TBG around the angle $\sim 1.13^{\circ}$. The geometry of the moiré pattern can be dynamically and reversibly switched between trigonal and square. The square pattern arises from a combination of twist and shear strain that minimizes the elastic energy.
Figure 14: Strain-induced triangular-rectangular moiré pattern evolution in twisted bilayer WSe2 with twist angle $\theta\approx 1.35^{\circ}$. (a) Schematic illustration of a piezoresponse force microscopy (PMF) technique to directly visualize moiré patterns on bent substrates. The substrate was mounted on molds with different curvatures to generate different strains. (b)-(f) PFM images of twisted bilayer WSe2 with substrate mounted on molds with (b) flat surface, strain magnitude $\varepsilon=0\%$, (c) with radius $R=25$ mm, $\varepsilon=0.25\%$, (d) with radius $R=16.7$ mm, $\varepsilon=0.38\%$, (e) with radius $R=13.9$ mm, $\varepsilon=0.45\%$, and (f) with radius $R=12.5$ mm, $\varepsilon=0.5\%$. Inserts are FFT images. The blue crosses indicate the moiré angle $\alpha$. The scale bars are 50 nm. Adapted under the terms of the $C$ 4.0 license from Ref. [ou2025continuous]. Copyright 2025 The author(s).

### v2

@@ III.6 Lattice relaxation
The geometrical description in the previous sections assumes that the lattices deform following the externally applied twist and strain. Real systems, however, tend to naturally relax in response to external perturbations, in order to minimize the energy. Relaxation effects in moiré heterostructures have been extensively studied in the literature [115, 85, 131, 18, 41, 79, 14, 66, 36, 54]. In the case of twisted bilayer graphene, the relaxation is particularly significant at very small (marginal) twist angles [115], whereby the energy-costly AA stacking regimes shrink and corrugate out of plane, leading to the formation of triangular AB/BA stacking separated by domain walls (DW) [131]. As a result, the relaxed moiré pattern of slightly twisted bilayer graphene can look markedly different than those expected for a rigid configuration. The relaxation acts, by itself, as a source of intrinsic inhomogeneous strain in the system [56, 53, 25].
@@
Similar relaxation effects are observed under external strain, although to date only few works have studied it. In the particular case of triangular moiré patterns induced by solely biaxial or shear strain (Section III.4.3), Cazeaux et al. reported also an expansion and contraction of AB/BA and AA stacking regimes, respectively, as the strain magnitude decreases [22]. A similar behavior was also recently reported by Kundu et al. in biaxially strained $\mathrm{MoS_{2}}$ bilayer [63]. There, the relaxed in-plane strain distribution is also modified by reducing and increasing, respectively, the strain magnitudes at the AB/BA and AA stacking regimes, which as in twisted systems, are separated by domain walls. Since these relaxation effects arise due to local minimization of the stacking and elastic energies, a similar behavior of DW formation is generally expected for non-triangular strained patterns. For instance, the quasi-unidimensional channels (Section III.4.1) also undergo a shrink and expansion of the AA and AB/BA stackings [22]. Another interesting relaxation effect is the formation of giant atomic swirls under pure biaxial heterostrain [82], whereby the soliton DW swirl around the AA stacking regimes [64].
@@
In general, the relaxed moiré configurations cannot be accounted by homogeneous twist and strain profiles, as considered in previous sections. Rather, they are characterized by a position dependent displacement vector at each atomic site, which minimizes the elastic and interlayer adhesion energies. The displacement vectors are typically obtained numerically [85, 18, 23], or analytically [114, 36, 54]. The geometrical description of relaxed moiré patterns arising from inhomogeneous strain is beyond the scope of this review.
III.7 Experimental determination of the strain
It is often the case that the strain in moiré heterostructures is not induced or controlled (Sec. IV), but rather arise naturally during the sample fabrication. In such situations, the strain configuration in the system is not known beforehand, so it can only be inferred by analyzing the geometrical and electronic properties of the moiré pattern. In this section, we briefly describe several methods to identify the strain and twist from the scanning tunneling microscopy (STM) topography of moiré structures [48, 6, 51, 60, 58, 127, 19].
A general formalism to identify the strain and twist from the atomically-resolved microscopy of any hexagonal moiré materials was first proposed by Artaud et al. [6]. In this method, a set of eight integers is identified by a Fourier analysis of STM images, from which one can then use the extended Wood’s notation to determine the geometry of the moiré superlattice. By using this method, Huder et al. [51] found that a small uniaxial heterostrain can suppress Dirac cones and induce flat bands in TBG. Kerelsky et al. [58] proposed another method to deal with STM images without atomic resolution, whereby by assuming that the samples suffer a uniaxial heterstrain one could extract the twist, strain magnitude and strain direction with the information of the moiré wavelengths only. Recently, Yu et al. [127] and Carrasco et al. [19] develop a simple minimization procedure that determines the twist and strain without presuming whether the strain is uniaxial or shear. By comparing the measured moiré wavelengths with the theoretical expressions for both strain types, the method identifies the twist-strain configuration that best matches the experimental results, providing a systematic way to determine the dominant strain component and the corresponding parameters directly from STM data.
All the above methods aim to infer the twist and strain configuration in the system through only the in-plane position of the atoms. The strain effect on the interlayer distance of the moiré pattern can be directly estimated from the STM topographic height profile [58].
@@ III.7 Experimental determination of the strain
Figure 10: The substrate bending technique. Schematic illustration of generating heterostrain in TBG on a flexible substrate (a) before and (b) after bending. R is the bending radius. (c) Calculated (red line) strain on the flexible substrate surface and experimental (black squares) uniaxial strain in bottom graphene. The thickness of the flexible substrate is 135 $\mathbf{\mu}$m. Inset shows the flexibility of the substrate with TBG device on top. Adapted under the terms of the $C$ BY license from Ref. [39]. Copyright (2021) The author(s).
@@ IV Experimental implementation of strain in moiré materials
Due to the nature of nanoscale dimensions and stability, strain is a powerful tuning knob to modify the geometrical and electronic properties of 2D systems [68, 71, 132, 108, 29, 70, 12, 75, 110, 26, 30, 20, 32, 69]. In particular, as we discuss in Sec. III, due to the presence of a moiré that acts as a magnifying glass, even very small strains can significantly modify the geometrical and electronic properties of the moiré patterns. In practice, small strain tend to be induced inevitably without control during the sample growth or fabrication, which we refer to as unintentional strain. Intentional strain may be also induced externally and designed carefully by well-established experimental techniques. In this section, we introduce three efficient experimental techniques developed to engineer strain in moiré materials. Then we present various strain-induced moiré pattern geometries generated in experiments.
Technique Strain range Moiré materials
Substrate out-of-plane bending up to $\sim 2.5\%$ TBG [39, 74], Graphene/hBN [55, 118], Twisted WSe2 [88].
Process-induced strain up to $\sim 3\%$ TBG [94], Twisted MoS2/WSe2 [134], Bilayer MoSe2 [33].
Sliding-based strain up to $\sim 0.81\%$ Graphene/hBN [99], TBG [19].
Table 1: Summary of strain implementation techniques, including the strain range and applicable moiré materials, discussed in Sec. IV.1.
@@ IV.1 Strain techniques
A large amount of experimental techniques have been introduced to generate strain in 2D materials. For example, strain can be generated by using different microactuators [96, 137], by applying hydrostatic pressure [5], by placing 2D materials on periodic structures constructed by nanopillars [71, 52], by stacking on piezoelectric stacks [74], or by bending a flexible substrate [118]. Here, we will mainly focus on several common approaches that are well-implemented in moiré materials. A brief summary of the key strain implementation techniques for engineering moiré materials is given in Table 1. Note that the strain techniques developed for moiré materials may not be compatible for moiré material-based devices. The strain approaches for 2D material-based devices are out of the scope of this review.
@@ IV.1.1 Substrate out-of-plane bending
Figure 11: Process-induced strain technique. (a) Schematic illustration of generating biaxial heterostrain by fully encapsulating a TBG with a highly stressed thin film. (b) Moiré patterns variation with biaxial heterostrain strength, where $\theta$ is the twist angle in the TBG. (c) Schematic illustration of generating uniaxial heterostrain by partially encapsulating TBG with a highly stressed thin film patterned into a stripe. (d) Moiré patterns variation with uniaxial heterostrain strength, and strain tunable Hubbard model parameters. The Hubbard model parameters $U$ and $t$ are the Coulomb repulsion and hopping terms, respectively. In the process-induced strain approach, the strain magnitude is directly proportional to the applied thin force on the stressor layer. Reprinted from Ref. [94], with the permission of AIP Publishing.
@@
Figure 12: Sliding-based strain technique. (a)-(c) Moiré patterns generated by hexagonal lattices with a lattice mismatch $\delta=10\%$ and (a) no strain, (b) $\varepsilon=2$% uniaxial heterostrain, (c) $\varepsilon=10$% uniaxial heterostrain. (d) Slide view schematic of a strained open-face graphene-hBN device. The graphene is stretched by pushing the right electrode with an AFM tip (gray triangle). Red lines show regions where the metal is strongly adhered through the O2–plasma treatment before deposition. (e) Optical image of a device. Red dashed regions are the O2–plasma treated electrode. Purple box labels the graphene channel. Scale bar is 4 $\mathbf{\mu}$m. (f)-(h) 100 nm $\times$ 100 nm CAFM images of graphene-hBN moiré patterns. The $x$-axis of the CAFM is along the graphene channel, parallel to the stretching direction. Insets show FFT of the images. Fitting the wavevectors results in the strain values $\varepsilon$ with strain angles $\phi$ relative to the +x+x-axis, and twist angle $\theta$ between graphene and hBN as (0.04%, +80 $\pm$ 20, -0.12), (0.34%, +1 $\pm$ 5, 0.02) and (0.04%, +17 $\pm$ 4, 0.02) for (f), (g) and (h), respectively. Reprinted with permission from Ref. [99]. Copyright 2024 American Chemical Society.
@@
One strain technique for moiré materials is based on in-plane or out-of-plane stretching or bending the underlying substrate or active 2D materials [118, 39, 55, 88]. In this technique, the active 2D layer can be suspended or directly stacked on the substrate. We refer to this method as substrate bending technique. A similar method to generate strain, through in-plane bending of one active 2D layer by the tip of an atomic force microscope (AFM), will be detailed in the following subsection [55].
@@
Figure 10 presents a generalized approach of dynamically tuning heterostrain in TBG via the substrate out-of-plane bending [39]. The setup consists of TBG assembled on a flexible polymethyl methacrylate (PMMA)-coated polyethylene terephthalate (PET) substrate. The twist angle between the two graphene layers is large (about $\theta\sim 13.2^{\circ}$ in the experiment) in order to minimize the interlayer friction force in TBG. As a result, the bottom graphene layer is separated from both the top graphene and the flexible substrate, and the strain generated from the substrate is only transferred to the bottom graphene. Upon bending the PET substrate, because the bottom graphene layer is strongly adhered to the substrate, it becomes uniaxially stretched. The top graphene experiences a negligible amount of deformation because it is bonded by the weak van der Waals force. Consequently, a uniaxial heterostrain is generated in the TBG moiré pattern. The strain strength can be dynamically and in situ modulated by changing the bending radius $R$, following the relation $\epsilon\approx t/2R$, where $t$ is the substrate thickness (Fig. 10(b)). The maximum strain strength achieved by out-of-plane bending in TBG with $\theta\sim 13.2^{\circ}$ was $\sim 1.3\%$ (Fig. 10(c)). The formation of heterostrain can be confirmed by the Raman spectroscopy, which presents triple G peaks with one degenerate mode G peak from unstrained graphene, and two split G+ and G- peaks from strained graphene.
@@ IV.1.2 Process-induced strain
Another common approach is a process-induced strain technique, which is based on the deposition of stressed thin films onto the active 2D materials [93, 94, 135, 134, 73]. In Ref. [94], Peña et al. utilized this method to generate designable layer-by-layer strains, on the device-level, with full control over the type of strain (uniaxial or biaxial, tensile or compressive) in moiré systems. By choosing optically transparent stressors, the intentionally induced strain could be characterized through Raman spectroscopy.
@@
Figure 13: Continuous tuning of twist angle and strain in a graphene–hBN heterostructure. (a) Schematic of the continuous tuning approach. The twist angle and the strain are manipulated by bending a 2D material ribbon through a nanomanipulator and the AFM tip. (b) Sketch of a distorted moiré pattern, magnifying the effect of strain in the low-angle limit. (c) (Left panel) AFM images of a bMLG on hBN before (top) and after (bottom) bending, (middle panel) PFM scans taken at the marked spots (white cross and dot) indicated in the bMLG and (right panel) their corresponding FFTs. From Ref. [55]. Reprinted with permission from AAAS.
@@ IV.1.3 Sliding-based strain
In Ref. [99], Sequeira et al. presented a sliding-based technique to manipulate the uniaxial heterostrain in graphene-hexagonal boron nitride (hBN) and study the strain effect on moiré geometry. To this end, they fabricated an open-faced graphene device on a hBN substrate, with a fixed source electrode and a movable drain electrode. In the electrode region, a metal handle is strongly adhered to the graphene layer (red line in Fig. 12(d)). Consequently, the drain electrode can be moved laterally by pushing it with an AFM tip, inducing heterostrain in the graphene layer. To selectively adhere the electrode to only the graphene layer, a light O2-plasma treatment is applied before depositing the Cr/Au handles. The electrodes are used to both selectively strain and measure transport (Fig. 12(e)). Importantly, the movable electrode has to have sufficient sliding friction to retain the strain in graphene when the AFM tip is retracted. As a result, an open-face graphene sample can be progressively strained, independently of the hBN substrate. The strain effect was measured by performing the conductive AFM (CAFM) measurement on the open graphene channels (Figs. 12(f)-(h)). As seen, the initial graphene aligned with hBN, forming a moiré pattern of length 14 nm, evolves into elongated moiré patterns by stretching the graphene with the AFM tip.
Recently, a similar sliding-based strain method was developed by Carrasco et al. [19]. During moiré sample growth, due to the different thermal expansion between graphene and the underlying substrate, graphene layers experience different sources of strain, resulting in graphene wrinkles in the graphene moiré sample. By mechanically manipulating the graphene wrinkles with a STM tip, they actively tune local strain in TBG around the angle $\sim 1.13^{\circ}$. The geometry of the moiré pattern can be dynamically and reversibly switched between trigonal and square. The square pattern arises from a combination of twist and shear strain that minimizes the elastic energy.
Figure 14: Strain-induced triangular-rectangular moiré pattern evolution in twisted bilayer WSe2 with twist angle $\theta\approx 1.35^{\circ}$. (a) Schematic illustration of a piezoresponse force microscopy (PMF) technique to directly visualize moiré patterns on bent substrates. The substrate was mounted on molds with different curvatures to generate different strains. (b)-(f) PFM images of twisted bilayer WSe2 with substrate mounted on molds with (b) flat surface, strain magnitude $\varepsilon=0\%$, (c) with radius $R=25$ mm, $\varepsilon=0.25\%$, (d) with radius $R=16.7$ mm, $\varepsilon=0.38\%$, (e) with radius $R=13.9$ mm, $\varepsilon=0.45\%$, and (f) with radius $R=12.5$ mm, $\varepsilon=0.5\%$. Inserts are FFT images. The blue crosses indicate the moiré angle $\alpha$. The scale bars are 50 nm. Adapted under the terms of the $C$ 4.0 license from Ref. [88]. Copyright 2025 The author(s).